OpenStudy (anonymous):

find the range of the function G: x --> x^2-4x+4; Domain= {0, 1, 2, 3}

OpenStudy (anonymous):

Because the domain is four distinct integers, and not any numbers in between, we must figure out the solution to each number in the domain to find the range. 0^2-4(0)+4=4 1^2-4(1)+4=1 2^2-4(2)+4=0 3^2-4(3)+4=1 The range is {0,1,4}, the only possible outputs.

OpenStudy (anonymous):

i love you